OpenStudy (anonymous):

Which of the following is equivalent to 2^4 =16? Logarithms

jimthompson5910 (jim_thompson5910):

hint: \[\large b^x = y\] turns into \[\large \log_{b}(y) = x\]

OpenStudy (anonymous):

would it be log2^16=4?

jimthompson5910 (jim_thompson5910):

correct, it is \[\large \log_{2}(16) = 4\]

OpenStudy (anonymous):

can you answer another one?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Which is equivalent to log 6^37 rounded to three decimal places?

jimthompson5910 (jim_thompson5910):

use the change of base rule \[\large \log_{b}(x) = \frac{\log(x)}{\log(b)}\]

OpenStudy (anonymous):

would it be 1.677?

jimthompson5910 (jim_thompson5910):

no that's incorrect

OpenStudy (anonymous):

oh is see its 2.015

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

thanks man

jimthompson5910 (jim_thompson5910):

np